OpenStudy (anonymous):

check please i pick the 3rd one Select the equations that are parallel and perpendicular to y = x + 5 and that pass through the point (-2, -1). parallel: y = -x - 1 perpendicular: y = x + 2 parallel: y = x - 1 perpendicular: y = -x + 1 parallel: y = x + 1 perpendicular: y = -x - 3 parallel: y = 2x - 2 perpendicular: y = -2x - 1

OpenStudy (anonymous):

perpendicular lines have slopes that are negatively reciprocal for this case the slope is 1. negative reciprocal of 1 is -1 With the general equation, y = mx + b. We know that the line has slope, -1 y = -x + b Now substitute x = -2 and y = -1 -1 = - (-2) + b -1 = 2 + b b = -3 Final equation is y = -x - 3

OpenStudy (anonymous):

and for parallel lines they have the same slope

OpenStudy (anonymous):

You are right the answer is the third one.